RHIC Run-11 Chronicle History
Jan. 8 RHIC Blue setup started. Seen first turn in Blue.
Jan. 9 Beam circulated in Blue with snake on. Closed orbit is close to R0 without hysteresis.
Jan. 10 Beam captured in Blue with 28MHz cavity and snake on. BPMs, WCM, Damper and Artus are all set up properly. Some polarimeter setup was done.
Jan. 11 Blue injection optics measured. BPM polarity checked. Feedback setup done with blue. Polarimeter setup made good progress.
Jan. 12 Problem with yellow injection kicker timing. Working on it around clock (three shifts).
Jan. 13 Smoking gun found: backup trigger (instead of master trigger) cable was in the timing module.Some polarimeter setup was done.
Jan. 14 Beam circulated in yellow with snake on.
Jan. 15 Beam captured in yellow with 28MHz cavity and snake on. BPMs, WCM, Damper and Artus are all set up properly.
Jan. 16 Power supply work is done during the day. Setup for ramp. LINAC mod8 problem is getting worse.
Jan. 17 First successful ramp after 4 tries.
Jan. 18 9MHz capture done. Seen glitches in quads on ramps at the same time.Chromaticity measured on the ramp and corrected.
Jan. 19 Maintenance day. Ramp development continues. Working on radial and chromaticity control on the ramp. Yellow polarimeter setup at injection.
Jan. 20 . Moved the separation bump from trim to main, namely, design orbit. 9MHz cavity commissioning.
Jan. 21 . Working on radial and chromaticity control on the ramp. First successful ramp with 9MHz cavity but lower intensity.
Jan. 22 .Reduced slam of gamma_t slam at t34. Adjust chromaticity on the ramp so that >90% transmission efficiency achieved. 28X28 ramp successful. Landau cavities are on.
Jan. 23 Successful 28X28 ramp. Beam is polarized at store. Continue to increase bunch number.
Jan. 24 Polarimeter setup. Successful 9MHz cavity ramps with Landau cavity on.
Jan. 25 28X28 store overnight. Polarization is about 40-50% at store. 9MHz cavity commissioning.
Jan. 26 Ring access for six hours to fix jet and many other jobs. Difficulty to provide 28X28 store overnight.
Jan. 27 Rotator ramp development. All polarimeter are functioning. Difficulty to provide 56X56 store overnight.
Jan. 28 9MHz cavity development.Rebucketed with 28MHz. Difficulty to provide 56X56 store overnight.
Jan. 29 Provide 56X56 store overnight. Working on clean up the ramp and chromaticity feedback development.
Jan. 30 Provide 56X56 store overnight. 28x28 ramp with rotator ramp was a success.
Jan. 31 Provide 56X56 store overnight. The ramp with gamma+tr quads on at a later time was a success, the tune glitch is largely gone. Pseudo-rotator ramp was commissioned.
Feb. 4 109x109 collisions for overnight store. 9MHz cavity commisioned. Collimator setup was retuned.
Feb. 5 109x109 collisions for overnight store. New orbit feedback algorithm commissioned but not finished.
Feb. 6 109x109 collisions for overnight store. New orbit feedback algorithm commissioning was done. Large blue emittance growth in overnight store.
Feb. 7 AGS RF station water leak and warm snake P/S problem prevented any beam development. Large blue emittance growth was still seen in overnight store.
Feb. 8 Found the source of blue emittance growth: blue abort kicker module B. Large blue emittance growth was gone in overnight store.
Feb. 9 Access for five hours. The noise from blue abort kicker module B came back. Another fix effort for eight hours. Overnight store lost due to yellow abort kicker prefire. A few ramps lost in no-rotator ramp.
Feb. 10 The noise from blue abort kicker module B came back again but it comes and goes. Develop ramp pp11v7(separate energy ramp and last beta squeeze). Provide overnight store with the blue noise presence.
Feb. 11 Spent the day to located the source of the blue dump kicker. And it is fixed. fixed a few loose end on the ramp setup (tune feedback in no-rotator ramp, automatic orbit correction, drift compensation). Provide 109x109 store overnight.
Feb. 12 Provide 109x109 stores.
Feb. 13 Provide 109x109 stores.
Feb. 14 Cryo lost power. Move maintenance day to today.
Feb. 15 Physics stores whole day.
Feb. 16 9MHz cavity setup did not go well due to LLRF problem. Then PASS system communication problem holded RHCI off by 5 hours. Troubleshooting RHIC LLRF consumed the next seven hours. No physics store at all.
Feb. 17 Rotator ramp development.
Feb. 18 Fix LLRF problem. Bump closure measurement. Physics store.
|